Find the constant a so that e^xy+y=x-1 is an implicit solution to dy/dx=(e^(-xy)+ay)/(e^(-xy)+x)
1
Expert's answer
2020-06-01T19:28:09-0400

We know that "e^{xy}+y = x-1" or "e^{xy}+y-x+1 = 0" . Let us take the derivative of the last expression with respect to x:

"e^{xy}\left( y+x\dfrac{dy}{dx}\right) + \dfrac{dy}{dx} - 1 + 0 = 0." Therefore, "\dfrac{dy}{dx} = \dfrac{1-ye^{xy}}{xe^{xy}+1} = \dfrac{e^{-xy}-y}{x+e^{-xy}}."

We can see that "e^{-xy}-y" should be equal to "e^{-xy}+ay" , therefore "a=-1."